06 March 2008

RHIC Running:  APEX ran yesterday from 9:00 - 16:00.  Measured spin tune at injection. 

There was a measurement of Yellow polarization at injection and store (fill 9988) that appeared to show improved polarization transmission on the ramp. 

For fill 9989, we filled 1.7/1.6e11 for 13kHz in PHENIX ZDC at store, with large Yellow vertical emittance initially.  Polarization went from 52/38 to 49/40 at store.  Yellow ramp efficiency is only 89%. 

PHENIX made a brief access to attempt to fix a Drift Chamber HV problem.  It was not successful. 

This morning PHENIX made another access to address the DC HV problem again and determined they need a longer ~1 hour access to replace a HV mainframe. 

At the end of the first store today there will be a test of the 10Hz oscillations by turning off the cryo flow.  Todd will coordinate these tests.

Due to problems with the AGS Corrector magnet, beam is down this morning, and both STAR(PP2PP) and PHENIX will be taking accesses of ~1 hour to address problems.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020